# Non-radial oscillations of hadronic neutron stars, quark stars, and hybrid stars : Estimation of $f$, $p$, and $g$ modes

### v1

@@ V.2 Quark Stars
We next study the structural and oscillation properties of the SQSs in Fig. 3. The mass-radius dependence of the QSs with different models are shown in Fig. 3a while corresponding tidal deformability in Fig. 3b. The parameters chosen for the v-Bag, NJL, and CFL models are according to Tabs. 2, 3, and 4, respectively. Our estimates of mass and radius for the v-Bag and NJL models are consistent with Sen and Guha (2022) and Lopes and Menezes (2021); Masuda et al. (2013), respectively. In accordance with other works in literature, we find that in case of v-Bag and NJL models, greater quark repulsion results in more massive SQSs with greater radius while in case of CFL SQSs higher value of gap parameter yields comparatively massive SQSs with larger radius. Out of all, although the models v-Bag 1, v-Bag 2, NJL 2, and CFL 2 satisfy all the constraints on the mass-radius plane, but the v-Bag 2 satisfies neither the constraint from GW170817 nor from GW190814.
Figs. 3c and 3d depict the variation of $f$ and $p$ mode non-radial oscillation frequencies of the SQSs, respectively. Comparing the two figures, we find that the $p$ mode is quite sensitive to the low density EoS than the $f$ mode. In absence of a bag constant, the EoS of the NJL model can be obtained from a much lower density compared to those of the v-Bag and CFL quark models. The latter models incorporate the bag constant in the EoS and in the expression of pressure, $B$ appears in a subtractive form Lopes et al. (2022); Lugones and Horvath (2002) that gives an unstable negative pressure range at low density. Thus we find from Fig. 3d that the $p_{1}$ mode frequency of the NJL model at low mass is quite low compared to that of the v-Bag and CFL models. In other words, the $p$ mode frequency is more dependent on the model considered. Similar to HNS, it can also be seen in case of QSs that $R_{1.4}$ is well correlated with $f_{f_{1.4}}$ because $R_{1.4}$ increases or $f_{f_{1.4}}$ decreases in the order CFL 1 $<$ NJL 1 $<$ CFL 2 $<$ NJL 2 $<$ v-Bag 1 $<$ v-Bag 2. This correlation is totally lost in case of $f_{{p_{1}}_{1.4}}$.
@@ V.3 Hybrid Stars
We proceed to display the estimates of $f$, $p_{1}$, and $g$ mode frequencies of the HSs in the Figs. 7 and 7 for the KIDS-A and KIDS-D hadronic models, respectively. Similar to HNSs and QSs, we observe that the negative correlation between $R_{1.4}$ and $f_{f_{1.4}}$ for HSs in Figs. 6a and 6a. Contrary to HNSs and QSs, it is interesting to note that the negative correlation between $R_{1.4}$ and $f_{{p_{1}}_{1.4}}$ exists for HSs in Figs. 6b and 6b. Unlike $f$ mode, the $p_{1}$ mode frequency of HSs is much greater than that of the HNSs even at high mass regime. However, due to different uncertainties pertaining to EoS based on different composition of compact star, it is still not possible to comment on whether $p$ mode can distinguish between HNS and QS. For the purpose, the $g$ mode frequency needs to be detected, which can support the possibility of phase transition and the existence of HSs. Theoretically, $g$ mode appears only in case of a discontinuity in density indicating a phase transition. In the present work, we obtain well defined $g$ mode frequencies corresponding to different HS configurations in Figs. 6c and 6c. We notice that the $g$ mode frequency decreases slightly with increasing mass and $f_{g_{1.4}}$ is negatively correlated to $R_{1.4}$ and $\Lambda_{1.4}$. Interestingly, there is substantial decrease in $g$ mode frequency with increasing values of $G_{V}$. Therefore, $g$ mode may also provide an idea of the strength of quark repulsion in HSs. For example, with the KIDS-D model, the change in the values of $f_{f_{1.4}}$, $f_{p_{1.4}}$, and $f_{g_{1.4}}$ for the maximum (0.37) and minimum (0.33) values of $G_{V}$ are 2.57%, 11.82%, and 20.86%, respectively. Similarly, with the KIDS-A model, the change in the values of $f_{f_{2.01}}$, $f_{p_{2.01}}$, and $f_{g_{2.01}}$ for the maximum and minimum values of $G_{V}$ are 0.8%, 2.32%, and 3.34%, respectively. Therefore, irrespective of mass of the HSs, $f_{g}$ is most affected by the strength of quark repulsion, followed by $f_{p}$.
@@
In Fig. 8 we compare the range of $f$, $p$, and $g$ mode frequencies of the HSs for example with KIDS-A model and $G_{V}$=0.35 fm2. We find that the magnitude of the three modes have three distinct range and in future the detection of these three modes can be well identified individually. Compared to $f$ and $p$ modes, the $g$ mode frequency is quite feeble and almost independent of mass.
The hadron-quark interface plays an important role in determining the properties of HSs. Therefore the properties at the transition are crucial. The estimates of the transition properties like the transition densities ($\rho_{t}^{H}$ and $\rho_{t}^{Q}$), transition mass ($M^{t}$), transition radius ($R^{t}$) and the $f$, $p_{1}$ and $g$ mode frequencies at transition ($f_{f}^{t}$, $f_{p1}^{t}$, and $f_{g}^{t}$) of HSs are tabulated below in Tab. 5.
@@
In Fig. 9 we study the variations of the eigenfunctions $W$ (Fig. 9a) and $V$ (Fig. 9b) with respect to the radius of 1.4 $M_{\odot}$ HS for example with KIDS-D model and $G_{V}$=0.35 fm2. The number of nodes corresponding to different modes of oscillation are visibly manifested in the figures. $p_{1}$ mode contains one node as both the eigenfunctions $W$ and $V$ become zero only once in the range $0<r<R$ for finite $r$, whereas $f$ mode does not have any node and $g$ mode shows phase transition. For $f$ mode, $W$ ($V$) increases (decreases) monotonically from the core to the surface of the star. We observe a glitch around $r/R$ = 0.7 indicating a phase transition. For $p$ mode, $W$ increases slightly from the core upto $r/R\approx$ 0.7 and then decreases sharply while $V$ decreases from the core slightly upto $r/R\approx$ 0.7 and then shows a stiff increase. For the $g$ mode, the variation of $W$ increases from the core upto $r/R\approx$ 0.7 and then falls drastically whereas $V$ decreases from the core upto $r/R\approx$ 0.7 and then shows a sudden jump and finally a slightly increasing (almost constant) variation towards the surface. This sudden jump in $V$ for $g$ mode is the most prominent indication of phase transition. This is because at the hadron-quark interface, the junction conditions given by Eqs. (15) and (16) indicate an abrupt change in $V$ while $W$ remains constant at the point of transition. Therefore the slightly decreasing part of $V$ from core to $r/R\approx$ 0.7 indicates the pure quark phase, followed by the vertical jump in $V$ indicating the jump in density due to the phase transition and finally the slightly increasing part of $V$ indicates the hadronic phase at higher radius.
@@
Fig. 11 is dedicated to study the universal relations in terms of $f_{f}$. For the purpose we consider only those models satisfying all the astrophysical constraints on the $M-R$ relations in Figs. 1a, 3a, 5a, and 5b and constraints on the $\Lambda$-$M$ plane from GW170817 and/or GW190814 in Figs. 1b, 3b, 5c, and 5d. In Fig. 11a we observe that angular frequency ($\omega_{f}$) corresponding to the $f$ mode, normalized with the mass, varies universally with the compactness $C$ for the different HNSs, QSs and HSs. This signifies that $f_{f}$ is almost independent of the composition of the star. Our linear fit for the universal relation reads as
@@
where, $a$ = 200.41 and $b$ = -5.65 are the fitting coefficients in the units of kHz-km with dimensionless $C$. Our fit is extremely close to the ones obtained by DKBP Das et al. (2021) and PC Pradhan and Chatterjee (2021), which are obtained for HNSs in presence of the hyperons using the relativistic Cowling approximation. This emphasizes that the $f$ mode frequency is nearly independent of the composition of the star and the underlying EoS. It is worth mentioning that this characteristic of the $f$ mode frequency is observed even with the calculations involving GR conditions Zhao and Lattimer (2022). It is also seen that the mass-scaled $f$ mode angular frequency of QSs is slightly less than that of HNSs and HSs.
In Fig. 11b we show that the $f$ mode frequency ($f_{f_{1.4}}$) of the 1.4 $M_{\odot}$ star bears well correlation with $\Lambda_{1.4}$ considering the different compositions and EoSs.
@@
We have already seen that the $f_{{p_{1}}_{1.4}}$ is not well-correlated with $R_{1.4}$ in case of HNSs and QSs although we find that for HSs $f_{{p_{1}}_{1.4}}$ is correlated with $R_{1.4}$ and $\Lambda_{1.4}$. Therefore it is not necessary to test the universality of $\omega_{p}M$ with respect to $C$ or of $f_{{p_{1}}_{1.4}}$ with respect to $R_{1.4}$ or $\Lambda_{1.4}$. However, $g$ mode is of special interest in this work. Therefore in Fig. 12 we check the universality of $g$ mode frequency. In Fig. 12a we show the variation of $\omega_{g}M$ with respect to $C$ for the HSs. Unlike the universal feature shown by $\omega_{f}M$ in Fig. 11a, we find $\omega_{g}M$ spans a much wider region (shaded) that widens more with increasing values of $C$ implying that unlike $f$ mode neither $p$ nor $g$ mode shows any universality in terms of the angular frequencies scaled with mass.
In the present work, phase transition is quite early with low transition mass (density) as seen from Tab. 5. Therefore we obtain the canonical stars (1.4 $M_{\odot}$) as hybrid ones. As we have already noticed there exists negative correlation between $f_{g_{1.4}}$ and $\Lambda_{1.4}$. This is displayed in Fig. 12b which also shows that the $f_{g_{1.4}}$-$\Lambda_{1.4}$ relationship (fitted) is also linear.
In the present work the oscillation properties are calculated with Cowling approximation. The quantitative results are supposed to deviate in the full GR calculations. However, it is well known that the qualitative results in both the treatments do not vary. This is supported by our correlation results in this work. For example, the correlation results in full GR treatment in Kunjipurayil et al. (2022) also ensure that irrespective of mass of HNSs, $f_{f}$ shows strongest correlation with $K_{\rm{sym}}$ while $f_{p_{1.4}}$ has strongest correlation with $L$ compared to other nuclear matter parameters. Moreover, correlation between $f_{f}$ and $K_{0}$ is found to be slightly decreasing with increasing mass of HNSs and in both Cowling approximation and full GR treatment the value is around 60% in Kunjipurayil et al. (2022). This value is very close to our estimation of $f_{f}$-$K_{0}$ as seen from Fig. 2. Similarly, the $f_{f}$-$L$ correlation decreases with increasing mass in full GR condition Kunjipurayil et al. (2022). In the present work we also find that $L$ is more correlated to $f_{f_{1.4}}$ than $f_{f_{2.01}}$ (Fig. 2). Also, it is already seen in Tonetto and Lugones (2020); Zhao et al. (2022) that Cowling approximation is reasonably suitable for calculation of $g$ mode frequency.
@@ VI Summary
In this work, we studied the non-radial oscillation frequencies arising from the quakes in the compact objects such as HNSs, SQSs, and HSs. $f$ and $p$ modes are examined for the HNSs, SQSs, and HSs. $g$ mode frequency is also calculated for HSs in addition to the $f$ and $p$ modes. EoS of the HNSs, SQSs, and HSs are strongly dependent on the model. Such a model dependence is thoroughly explored by adopting six models for hadronic matter, three models with varying parameter values for the strange quark matter. For the HSs we consider two hadronic models with three values of repulsion strength of the pure quark matter.
@@
For HNSs, EoSs of the chosen RMF models are substantially stiffer than the KIDS model. As a consequence the mass, radius, and tidal deformability of the HNSs are clearly distinguished between the KIDS and the RMF models. The $f$ mode frequency of the 1.4 $M_{\odot}$ star is strongly correlated with $R_{1.4}$. $f_{f}$ increases exactly in the reverse order of the radius. So we have the variation of $f_{f_{1.4}}$ in the order GM1 $<$ DD2 $<$ KIDS-C $<$ KIDS-A $<$ KIDS-D $<$ KIDS-B. In this work the frequency of the softer hadronic EoS resides in the range (2.2 - 2.3) kHz while the stiffer ones in (2.0 - 2.05) kHz. So accurate measurement of $f_{f_{1.4}}$ can put a stringent constraint on the stiffness of EoS. Results of $f_{p_{1.4}}$, unlike $f_{f_{1.4}}$, show no systematic correlation with the bulk properties of HNSs. However, it is arranged exactly in the reverse order of the slope parameter $L$ of the symmetry energy. In the KIDS model we have $f_{p_{1.4}}$approx 6.15, 6.4, and 6.8 kHz with the KIDS-A, C, and D models, respectively. The interval between the models is large enough that if $f_{p_{1.4}}$ is measured accurately, it could provide exclusive constraints to reduce the uncertainty of $L$.
@@
For the SQSs the bulk properties of the stars depend more strongly on the model. Excluding the models that have maximum mass below 2 $M_{\odot}$, $f_{p_{1.4}}$ is in the range (2 - 2.2) kHz. So they are in similar range as that of HNSs. For the $p$ mode, however, only the NJL 2 model gives $f_{p_{1.4}}$ = 5.5 kHz which is similar to that of HNSs, but the other models collectively predict $f_{p_{1.4}}$ in the range (7.5 - 11) kHz. This range is completely separated from that of $f_{f_{1.4}}$. So the measurement of $p$ mode could be useful to identify the existence of SQSs.
@@
In the calculation of the HSs, in addition to the $f$ and $p$ modes, we consider the $g$ mode that originated from the first order phase transition in the core. The effect of the symmetry energy is accounted by adopting KIDS-A and KIDS-D models and the uncertainty in the quark matter EoS is incorporated with the strength of the vector repulsion in the v-Bag model. $f_{f_{1.4}}$ of HSs is in the range (2.25 - 2.35) kHz, which overlaps with that of HNSs. So $f$ mode is not suitable to distinguish between HNSs and HSs. $p$ mode results vary opposite to the $f$ mode. So $f_{p_{1.4}}$ of HSs are clearly distinguished from those of the HNSs, and the dependence of the symmetry and $G_{V}$ also appears clearly. Thus the measurement of $p$ mode is expected to provide diverse and rich story about the various aspects of the dense matter theory and properties of compact objects. $g$ mode shows more dramatic dependence on the symmetry energy. For the KIDS-A model, $f_{g_{1.4}}$ is in the range (0.77 - 0.81) kHz but it is (0.57 - 0.72) kHz for the KIDS-D model. So there is no overlap between the KIDS-A and KIDS-D models. The dependence on $G_{V}$ is more prominent in the the KIDS-D model. With $G_{V}$ = 0.33 fm2, $f_{g_{1.4}}$approx 0.72 kHz but it drops to 0.57 kHz when $G_{V}$ = 0.37 fm2. So $g$ mode can provide a unique constraint to reduce the uncertainty in the deconfined quark matter EoS in HSs in terms of quark repulsive strength.
@@
The relation between compactness $C$ and the normalized $f$ mode frequency $\omega_{f}M$ shows a universal behavior independent of mass, radius, and composition of the compact objects. Measurement of $f$ mode will thus provide a critical test for the general relativistic description of the compact objects and their properties. We also find strong correlations between the tidal deformability and the $f$ and $g$ mode frequencies of 1.4 $M_{\odot}$ star, independent of composition. So they complement each other to sharpen our understanding of the dense matter properties.
@@
Our analysis from the correlation study suggests that disregarding the mass of HNSs $f_{f}$ is strongly correlated to $K_{\rm{sym}}$, radius, and tidal deformability, moderately correlated to $K_{0}$ and $L$ and poorly correlated to $J$. $f_{p}$ shows strongest correlation with $L$ and moderate correlation with radius and tidal deformability. In case of SQS, irrespective of the mass, $f_{f}$ is very strongly correlated to both radius and tidal deformability while $p$ mode frequency for SQSs shows moderate correlation with $\Lambda$ but poor correlation with radius. For HSs, regardless of the mass, both $f_{f}$ and $f_{p}$ are well correlated to radius. However, only $f_{f}$ is well correlated to $\Lambda$, regardless of the mass of the HSs. $f_{g}$ also shows moderate correlation with radius, irrespective to the mass of HSs.
@@
Overall, the non-radial oscillations of the compact objects in the $f$, $p$, and $g$ modes have their own unique and independent characteristics. Measurements of the frequencies are promising to deepen and sharpen our understanding of the strongly interacting infinite matter at densities above the nuclear saturation. In the present work we have adopted Cowling approximation since the qualitative results of the $f$ and $p$ mode frequencies and the quantitative estimation of $f_{g}$ remain almost same in full GR treatment. However, it will be interesting to study the oscillation properties of the HNSs, SQSs, and HSs in full GR scenario in a near future work.

### v2

@@ V.2 Quark Stars
Next, we study the structural and oscillation properties of the SQSs in Fig. 3. The mass-radius dependence of the QSs with different models is shown in Fig. 3a while corresponding tidal deformability in Fig. 3b. The parameters chosen for the v-Bag, NJL, and CFL models are according to Tabs. 2, 3, and 4, respectively. Our estimates of mass and radius for the v-Bag and NJL models are consistent with Sen and Guha (2022) and Lopes and Menezes (2021); Masuda et al. (2013), respectively. In accordance with other works in literature, we find that in the case of v-Bag and NJL models, greater quark repulsion results in more massive SQSs with greater radius while in the case of CFL SQSs higher value of gap parameter yields comparatively massive SQSs with larger radius. Out of all the models, although the v-Bag 1, v-Bag 2, NJL 2, and CFL 2 satisfy all the constraints on the mass-radius plane, the v-Bag 2 satisfies neither the constraint from GW170817 nor from GW190814.
Figs. 3c and 3d depict the variation of $f$ and $p$ mode non-radial oscillation frequencies of the SQSs, respectively. Comparing the two figures, we find that the $p$ mode is quite sensitive to the low density EoS than the $f$ mode. In absence of a bag constant, the EoS of the NJL model can be obtained from a much lower density compared to those of the v-Bag and CFL quark models. The latter models incorporate the bag constant in the EoS and in the expression of pressure, $B$ appears in a subtractive form Lopes et al. (2022); Lugones and Horvath (2002) that gives an unstable negative pressure range at low density. Thus, we find from Fig. 3d that the $p_{1}$ mode frequency of the NJL model at low mass is quite low compared to that of the v-Bag and CFL models. In other words, the $p$ mode frequency is more dependent on the model considered. Similar to HNS, it can also be seen in the case of QSs that $R_{1.4}$ is well correlated with $f_{f_{1.4}}$ because $R_{1.4}$ increases or $f_{f_{1.4}}$ decreases in the order CFL 1 $<$ NJL 1 $<$ CFL 2 $<$ NJL 2 $<$ v-Bag 1 $<$ v-Bag 2. This correlation is totally lost in case of $f_{{p_{1}}_{1.4}}$.
@@ V.3 Hybrid Stars
We proceed to display the estimates of $f$, $p_{1}$, and $g$ mode frequencies of the HSs in the Figs. 7 and 7 for the KIDS-A and KIDS-D hadronic models, respectively. Similar to HNSs and QSs, we observe that the negative correlation between $R_{1.4}$ and $f_{f_{1.4}}$ for HSs in Figs. 6a and 6a. Contrary to HNSs and QSs, it is interesting to note that the negative correlation between $R_{1.4}$ and $f_{{p_{1}}_{1.4}}$ exists for HSs in Figs. 6b and 6b. Unlike $f$ mode, the $p_{1}$ mode frequency of HSs is much greater than that of the HNSs even at the high mass regime. However, due to different uncertainties pertaining to EoS based on different composition of compact star, it is still not possible to comment on whether $p$ mode can distinguish between HNS and QS. For the purpose, the $g$ mode frequency needs to be detected, which can support the possibility of phase transition and the existence of HSs. However, its detection is quite challenging in the recent future due its feeble tidal coupling. Theoretically, $g$ mode appears only in the case of a discontinuity in density indicating a phase transition. In the present work, we obtain well-defined $g$ mode frequencies corresponding to different HS configurations in Figs. 6c and 6c. We notice that the $g$ mode frequency decreases slightly with increasing mass and $f_{g_{1.4}}$ is negatively correlated to $R_{1.4}$ and $\Lambda_{1.4}$. Interestingly, there is substantial decrease in $g$ mode frequency with increasing values of $G_{V}$. Therefore, $g$ mode may also provide an idea of the strength of quark repulsion in HSs. For example, with the KIDS-D model, the change in the values of $f_{f_{1.4}}$, $f_{p_{1.4}}$, and $f_{g_{1.4}}$ for the maximum (0.37) and minimum (0.33) values of $G_{V}$ are 2.57%, 11.82%, and 20.86%, respectively. Similarly, with the KIDS-A model, the changes in the values of $f_{f_{2.01}}$, $f_{p_{2.01}}$, and $f_{g_{2.01}}$ for the maximum and minimum values of $G_{V}$ are 0.8%, 2.32%, and 3.34%, respectively. Therefore, irrespective of mass of the HSs, $f_{g}$ is most affected by the strength of quark repulsion, followed by $f_{p}$.
@@
In Fig. 8 we compare the range of $f$, $p$, and $g$ mode frequencies of the HSs for example with the KIDS-A model and $G_{V}$=0.35 fm2. We find that the magnitudes of the three modes have three distinct ranges. In future the detection of these three modes can be well identified individually. Compared to $f$ and $p$ modes, the $g$ mode frequency is quite feeble and almost independent of mass.
The hadron-quark interface plays an important role in determining the properties of HSs. Therefore, the properties at the transition are crucial. The estimates of the transition properties such as the transition densities ($\rho_{t}^{H}$ and $\rho_{t}^{Q}$), transition mass ($M^{t}$), transition radius ($R^{t}$) and the $f$, $p_{1}$ and $g$ mode frequencies at transition ($f_{f}^{t}$, $f_{p1}^{t}$, and $f_{g}^{t}$) of HSs are tabulated below in Tab. 5.
@@
In Fig. 9 we study the variations of the eigenfunctions $W$ (Fig. 9a) and $V$ (Fig. 9b) with respect to the radius of 1.4 $M_{\odot}$ HS for example with KIDS-D model and $G_{V}$=0.35 fm2. The number of nodes corresponding to different modes of oscillation is visibly manifested in the figures. $p_{1}$ mode contains one node as both the eigenfunctions $W$ and $V$ become zero only once in the range $0<r<R$ for finite $r$, whereas $f$ mode does not have any node and $g$ mode shows phase transition. For $f$ mode, $W$ ($V$) increases (decreases) monotonically from the core to the surface of the star. We observe a glitch around $r/R$ = 0.7 indicating a phase transition. For the $p$ mode, $W$ increases slightly from the core upto $r/R\approx$ 0.7 and then decreases sharply while $V$ decreases from the core slightly up to $r/R\approx$ 0.7 and then shows a stiff increase. For the $g$ mode, the variation of $W$ increases from the core up to $r/R\approx$ 0.7 and then falls drastically whereas $V$ decreases from the core up to $r/R\approx$ 0.7 and then shows a sudden jump and finally a slightly increasing (almost constant) variation towards the surface. This sudden jump in $V$ for the $g$ mode is the most prominent indication of phase transition. This is because at the hadron-quark interface, the junction conditions given by Eqs. (16) and (17) indicate an abrupt change in $V$ while $W$ remains constant at the point of transition. Therefore, the slightly decreasing part of $V$ from core to $r/R\approx$ 0.7 indicates the pure quark phase, followed by the vertical jump in $V$ indicating the jump in density due to the phase transition and finally the slightly increasing part of $V$ indicates the hadronic phase at higher radius.
@@
Fig. 11 is dedicated to the study of universal relations in terms of $f_{f}$. For this purpose, we consider only those models that satisfy all the astrophysical constraints on the $M-R$ relations in Figs. 1a, 3a, 5a, and 5b and constraints on the $\Lambda$-$M$ plane from GW170817 and/or GW190814 in Figs. 1b, 3b, 5c, and 5d. In Fig. 11a we observe that the angular frequency ($\omega_{f}$) corresponding to the $f$ mode, normalized with the mass, varies universally with the compactness $C$ for the different HNSs, QSs and HSs. This signifies that $f_{f}$ is almost independent of the composition of the star. Our linear fit for the universal relation is as follows.
@@
where, $a$ = 200.41 and $b$ = -5.65 are the fitting coefficients in the units of kHz-km with dimensionless $C$. Our fit is extremely close to the ones obtained by DKBP Das et al. (2021) and PC Pradhan and Chatterjee (2021), which are obtained for HNSs in the presence of the hyperons using the relativistic Cowling approximation. This emphasizes that the $f$ mode frequency is nearly independent of the composition of the star and the underlying EoS. It is worth mentioning that this characteristic of the $f$ mode frequency is observed even with the calculations involving GR conditions Zhao and Lattimer (2022). It is also seen that the mass-scaled $f$ mode angular frequency of QSs is slightly less than that of HNSs and HSs.
In Fig. 11b we show that the $f$ mode frequency ($f_{f_{1.4}}$) of the 1.4 $M_{\odot}$ star bears good correlation with $\Lambda_{1.4}$ considering the different compositions and EoSs.
@@
We have already seen that the $f_{{p_{1}}_{1.4}}$ is not well-correlated with $R_{1.4}$ in case of HNSs and QSs although we find that for HSs $f_{{p_{1}}_{1.4}}$ is correlated with $R_{1.4}$ and $\Lambda_{1.4}$. Therefore, it is not necessary to test the universality of $\omega_{p}M$ with respect to $C$ or of $f_{{p_{1}}_{1.4}}$ with respect to $R_{1.4}$ or $\Lambda_{1.4}$. However, $g$ mode is of special interest in this work. Therefore in Fig. 12 we check the universality of $g$ mode frequency. In Fig. 12a we show the variation of $\omega_{g}M$ with respect to $C$ for the HSs. Unlike the universal feature shown by $\omega_{f}M$ in Fig. 11a, we find $\omega_{g}M$ spans a much wider region (shaded) that widens more with increasing values of $C$. The result implies that unlike $f$ mode neither $p$ nor $g$ mode shows any universality in terms of the angular frequencies scaled with mass.
In the present work, phase transition is quite early with low transition mass (density), as seen from Tab. 5. Therefore, we obtain the canonical stars (1.4 $M_{\odot}$) as hybrid ones. We have already noticed that there exists a negative correlation between $f_{g_{1.4}}$ and $\Lambda_{1.4}$. This is displayed in Fig. 12b which also shows that the $f_{g_{1.4}}$-$\Lambda_{1.4}$ relationship (fitted) is also linear.
In this work the oscillation properties are calculated with Cowling approximation. The quantitative results are supposed to deviate from the full GR calculations. However, it is well known that the qualitative results from both the treatments do not vary. This is supported by our correlation results in this work. For example, the correlation results in full GR treatment in Kunjipurayil et al. (2022) also ensure that irrespective of mass of the HNSs, $f_{f}$ shows strongest correlation with $K_{\rm{sym}}$ while $f_{p_{1.4}}$ has the strongest correlation with $L$ compared to other the nuclear matter parameters. Moreover, the correlation between $f_{f}$ and $K_{0}$ is found to decrease slightly with increasing mass of HNSs. In both Cowling approximation and full GR treatment the value is around 60% in Kunjipurayil et al. (2022). This value is very close to our estimation of $f_{f}$-$K_{0}$ as seen from Fig. 2. Similarly, the $f_{f}$-$L$ correlation decreases with increasing mass in the full GR condition Kunjipurayil et al. (2022). In the present work, we also find that $L$ is more correlated to $f_{f_{1.4}}$ than $f_{f_{2.01}}$ (Fig. 2). Also, it is already seen in Tonetto and Lugones (2020); Zhao et al. (2022) the Cowling approximation is reasonably suitable for the calculation of $g$ mode frequency.
@@ VI Summary
In this work, we studied the non-radial oscillation frequencies arising from the quakes in the compact objects such as HNSs, SQSs, and HSs. $f$ and $p$ modes are examined for the HNSs, SQSs, and HSs. $g$ mode frequency is also calculated for the HSs in addition to the $f$ and $p$ modes. The EoS of the HNSs, SQSs, and HSs are strongly dependent on the model. Such a model dependence is thoroughly explored by adopting six models for hadronic matter, three models with varying parameter values for the strange quark matter. For the HSs we consider two hadronic models with three values of repulsion strength of the pure quark matter.
@@
For HNSs, the EoSs of the chosen RMF models are substantially stiffer than the KIDS model. As a consequence, the mass, radius, and tidal deformability of the HNSs are clearly distinguished between the KIDS and the RMF models. The $f$ mode frequency of the 1.4 $M_{\odot}$ star is strongly correlated with $R_{1.4}$. $f_{f}$ increases exactly in the reverse order of the radius. So we have the variation of $f_{f_{1.4}}$ in the order GM1 $<$ DD2 $<$ KIDS-C $<$ KIDS-A $<$ KIDS-D $<$ KIDS-B. In this work the frequency of the softer hadronic EoS resides in the range (2.2 - 2.3) kHz while the stiffer ones in (2.0 - 2.05) kHz. So, accurate measurement of $f_{f_{1.4}}$ can put a stringent constraint on the stiffness of EoS. The results of $f_{p_{1.4}}$, unlike $f_{f_{1.4}}$, do not show any systematic correlation with the bulk properties of HNSs. However, it is arranged exactly in the reverse order of the slope parameter $L$ of the symmetry energy. In the KIDS model we have $f_{p_{1.4}}$approx 6.15, 6.4, and 6.8 kHz with the KIDS-A, C, and D models, respectively. The interval between the models is large enough that if $f_{p_{1.4}}$ is measured accurately, it could provide exclusive constraints to reduce the uncertainty of $L$.
@@
For the SQSs the bulk properties of the stars depend more strongly on the model. Excluding the models that produce maximum mass below 2 $M_{\odot}$, $f_{f_{1.4}}$ is in the range (2 - 2.2) kHz. So, they are in range similar to that of HNSs. For the $p$ mode, however, only the NJL 2 model gives $f_{p_{1.4}}$ = 5.5 kHz which is similar to that of HNSs, but the other models collectively predict $f_{p_{1.4}}$ in the range (7.5 - 11) kHz. This range is completely separate from that of $f_{f_{1.4}}$. So the measurement of $p$ mode could be useful in identifying the existence of SQSs.
@@
In the calculation of the HSs, in addition to the $f$ and $p$ modes, we consider the $g$ mode that originated from the first order phase transition in the core. The effect of the symmetry energy is accounted for by adopting KIDS-A and KIDS-D models and the uncertainty in the quark matter EoS is incorporated with the strength of the vector repulsion in the v-Bag model. $f_{f_{1.4}}$ of HSs is in the range (2.25 - 2.35) kHz, which overlaps with that of HNSs. So, $f$ mode is not suitable to distinguish between HNSs and HSs. $p$ mode results vary opposite to the $f$ mode. So $f_{p_{1.4}}$ of HSs are clearly distinguished from those of the HNSs, and the dependence of the symmetry and $G_{V}$ also appears clearly. Thus, the measurement of $p$ mode is expected to provide a diverse and rich story about the various aspects of the theory of dense matter and the properties of the compact objects. $g$ mode shows more dramatic dependence on the symmetry energy. For the KIDS-A model, $f_{g_{1.4}}$ is in the range (0.77 - 0.81) kHz but it is (0.57 - 0.72) kHz for the KIDS-D model. So, there is no overlap between the KIDS-A and KIDS-D models. The dependence on $G_{V}$ is more prominent in the KIDS-D model. With $G_{V}$ = 0.33 fm2, $f_{g_{1.4}}$approx 0.72 kHz but it drops to 0.57 kHz when $G_{V}$ = 0.37 fm2. So $g$ mode can provide a unique constraint to reduce the uncertainty in the deconfined quark matter EoS in HSs in terms of quark repulsive strength.
@@
The relation between compactness $C$ and the normalized $f$ mode frequency $\omega_{f}M$ shows a universal behavior independent of the mass, radius, and composition of the compact objects. Measurement of $f$ mode will thus provide a critical test for the general relativistic description of the compact objects and their properties. We also find strong correlations between the tidal deformability and the $f$ and $g$ mode frequencies of the 1.4 $M_{\odot}$ star, independent of composition. So, they complement each other to sharpen our understanding of the dense matter properties.
@@
Our analysis of the correlation study suggests that disregarding the mass of HNSs $f_{f}$ is strongly correlated to $K_{\rm{sym}}$, radius, and tidal deformability, moderately correlated to $K_{0}$ and $L$ and poorly correlated to $J$. $f_{p}$ shows strongest correlation with $L$ and moderate correlation with radius and tidal deformability. In the case of SQS, irrespective of the mass, $f_{f}$ is very strongly correlated to both the radius and the tidal deformability, while the $p$ mode frequency for SQSs shows a moderate correlation with $\Lambda$ but a poor correlation with radius. For HSs, regardless of the mass, both $f_{f}$ and $f_{p}$ are well correlated to radius. However, only $f_{f}$ is well correlated to $\Lambda$, regardless of the mass of the HSs. $f_{g}$ also shows moderate correlation with radius, irrespective of the mass of HSs.
@@
Overall, the non-radial oscillations of the compact objects in the $f$, $p$, and $g$ modes have their own unique and independent characteristics. Measurements of the frequencies are promising to deepen and sharpen our understanding of the strongly interacting infinite matter at densities above the nuclear saturation. From the perspective of the upcoming detectors, the detection of $f_{f}$ is most plausible while $f_{g}$ is most challenging. In the present work we have adopted the Cowling approximation since the qualitative results of the $f$ and the $p$ mode frequencies and the quantitative estimation of $f_{g}$ remain almost the same in full GR treatment. However, it will be interesting to study the oscillation properties of the HNSs, SQSs, and HSs in full GR scenario in a near future work.